OpenStudy (anonymous):

The graph of g(x) is f(x) translated to the right 8 units and up 2 units. What is the function rule for g(x) given f(x)=xsquared?

OpenStudy (anonymous):

When you raise a graph k units, then the graph of y = x^2 becomes y = x^2 + k When you move (translate) the graph of y = x^2, say 5 units to the right, its new graph is y = (x-5)^2 Use the above to answer your question.

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

y=(x+8)^2